The sum of 2 numbers is 40. When the greater number is divided by the smaller, the quotient is 4 and the remainder is 5. Find the numbers.
Answer by xdragonfight(116)

first number=x 2nd number= y
The sum of 2 numbers is 40
so x+y=40
x/y = 4+5/y
x = 4y+5
x+y=40
4y+5+y=40
5y+5=40
5y=35
y=7
x=4y+5=4(7)+5=28+5=33
Answers: numbers are 33 and 7